OpenStudy (anonymous):

HELP PLEASE! Find the points of inflection of the graph of the function. (If an answer does not exist, enter DNE.) f(x) = 5 sin x + 5 cos x, [0, 2π] (x, y) = (smaller x-value) (x, y) = (larger x-value) Describe the concavity. (Enter your answers using interval notation.) concave upward : concave downward :

OpenStudy (anonymous):

Well you know what gives us points of inflection correct?

OpenStudy (anonymous):

no i don't @hilbertboy96

OpenStudy (anonymous):

ik i have to take the second derivative which would be -5 sin(x)-5cos(x), and set that to zero correct? that would be -5 sin(x)- 5cos(x)=0 then i have to factor out the 5? -5(sin(x)+cos(x))=0, but then i forget what to do next

OpenStudy (paxpolaris):

\[ -5 \sin(x)- 5\cos(x)=0\]\[ \implies \sin(x)+\cos(x)=0\].... next can you rewrite this sum of two sinusoidal function as a single sine function?

OpenStudy (anonymous):

sin(x)=0 ?

OpenStudy (paxpolaris):

sorry nvm what i said earlier ... \[ \sin(x)+\cos(x)=0\]\[\implies \sin(x)=-\cos(x)\]\[\implies \tan(x)=-1\]

OpenStudy (anonymous):

ok, so do i plug the -1 into the original function?

OpenStudy (anonymous):

@PaxPolaris

OpenStudy (paxpolaris):

no you need to solve for x still ...

OpenStudy (anonymous):

for x from the original function or the one above?

OpenStudy (paxpolaris):

in the domain: \(0\le x \le 2\pi\)\[\tan (x)=-1\]\[\implies x={3\pi \over4}\text { OR } x={7\pi \over4}\]

OpenStudy (anonymous):

ok, so i would use those and substitute that for x in the original function?

OpenStudy (paxpolaris):

right, to get y for the 2 inflection points

OpenStudy (anonymous):

ok, thanks